[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Thu Mar 30, 2017

TAFB Shift Log nhc.tsbadmin at noaa.gov
Thu Mar 30 18:35:26 UTC 2017


TAFB Shift Log for Thu Mar 30 2017 14:35:20 GMT-0400 (EDT)

Shift: Day Shift

Pacific Forecaster: C McElroy

Atlantic Forecaster: N Ramos

Surface Analyst: M Formosa

Technical Issues Summary:

GFE/NCP issues:

N-AWIPS issues: 1335L - I have worked with Paul on the upgrade to the  
Starfish NAWIPS workstation. Paul fixed a navigation problem involving the  
cursor going side to side but not up or down between monitors. Paul and  
Brian are working on observations not coming into NWX. SPFs and NTRANS were  
working on Starfish. CAM 1355L - Chris and Brian are upgrading monitors on  
Sheepshead and Starfish. CAM 1400L - Obs are now coming into NWX on  
Starfish. CAM

Other issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170330/f404da32/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list